### v1

@@ 2.6 Model Selection
According to Watanabe and Opper (2010), a statistical model is said to be regular if the map taking parameters to probability distributions is one-to-one and if its Fisher information matrix is positive definite. If a model is not regular, then it is said to be singular. If a statistical model contains a hierarchical structure or latent variables then the model is generally singular. In singular statistical models, the maximum likelihood estimator does not satisfy asymptotic normality. Consequently, standard model selection criteria like the Akaike Information Criterion (AIC) and the Bayes Information Criterion (BIC) are not always appropriate for model selection. WAIC (Widely Applicable Information Criterion) (Watanabe and Opper, 2010) can be used for estimating the predictive loss of Bayesian models, using a sample from the full-data posterior, and it is applicable to non-regular models such as the PM.
In the present paper, WAIC is calculated from Equations 5, 12, and 13 in Gelman et al (2014), and it is the log pointwise predictive density minus a correction for effective number of parameters to adjust for overfitting. According to Millar (2018), the marginalized WAIC might be more accurate for choosing the correct model.
We fit the model in a simulated data scenario to verify the number of times the correct model is selected by the WAIC conditional on all of the parameters (WAICc) and marginalized over $\bm{\tau}$ (WAICm). We generated 100 random membership vectors from a Dirichlet($\bm{\delta}$) distribution with shape parameter $\bm{\delta}$ = (0.5, 0.5, 0.5, 0.5). Using these membership scores, we simulated 100 partial membership models with $N=200$, $J=25$ and $K=4$ to match the football players application scenario. We ran the MCMC algorithm for 20000 iterations, keeping every 50th draw. We discarded the first 5000 iterations as burn-in. MCMC was run the model over a range of values of $K=1,...,8$. To assess convergence, we examined trace plots.
The simulation results are illustrated in Table 2 and the simulation confirms that the marginalized WAIC, is more accurate for choosing the right model, with $99\%$ of success.
Table 2: Number of times each $K$ corresponds to the minimum WAIC conditional to all the parameters (WAICc) and marginalized for $\bm{\tau}$ (WAICm). $N=200$, $J=25$, Number of runs=100, true $K=4$
WAICc WAICm
@@ 3 Serie A Football Players
We analyze the performance statistics of the 192 Serie A football players who played more than 1350 minutes during the 2022/2023 football season [^1] The data are freely available at https://fbref.com.. The analysis considers a set of 22 count variables for each player across the games that they played, selected to encompass the essential skills associated with each player’s role on the field.
@@
The performance variables included in the analysis and modelling are given as follows.
@@
  • PrgP – Progressive Passes: progressive Passes completed passes that move the ball towards the opponent’s goal line at least 10 yards from its furthest point in the last six passes, or any completed pass into the penalty area. Excludes passes from the defending 40% of the pitch.
@@
  • SoT – Shots on Target: does not include penalty kicks.
@@
  • PiFT – Passes into Final Third: completed passes that enter the 1/3 of the pitch closest to the goal, not including set pieces.
  • PPA – Passes into Penalty Area: completed passes into the 18-yard box, not including set pieces.
@@
  • PassLive – SCA (PassLive): completed live-ball passes that lead to a shot attempt.
  • PassDead – SCA (PassDead): completed dead-ball passes that lead to a shot attempt. Includes free kicks, corner kicks, kick offs, throw-ins and goal kicks.
@@
  • Err – Errors: Mistakes leading to an opponent’s shot
Partial membership (PM), mixed membership (MM) and mixture models are applied to the dataset, with the WAIC suggesting that the $K=4$, $K=5$ and $K=6$ profile models are optimal, as illustrated in Table 3.
Table 3: WAIC values for the partial membership, mixed membership and mixture models, with $K$ ranging between 2 and 8. The optimal model for each model type is highlighted in boldface.
WAIC $K=2$ $K=3$ $K=4$ $K=5$
PM 52237.45 44274.20 37692.89 47919.76
MM 35330.92 29833.80 28309.35 27898.17
Mixture 35385.72 29846.14 28369.55 27981.13
@@ 3.1 Partial Membership Model Results
Table 4 and Figure 2 give the mean profiles for the partial membership model with $K=4$.
@@
Table 4: PM model expected profiles means
Gls Ast PrgC PrgP Sh SoT KP PiFT
1 0.00 0.04 0.13 0.32 0.00 0.00 0.13 8.02
2 0.65 0.19 6.42 52.34 8.39 2.11 2.21 45.67
3 2.74 3.75 71.57 170.20 29.29 9.25 45.91 136.88
4 15.14 3.70 42.25 19.69 88.57 34.20 26.75 6.75
PPA CrsPA SCA PassLive PassDead TO ShToSh Def
1 0.19 0.11 1.45 1.09 0.77 0.01 0.00 0.07
2 1.79 0.36 8.18 7.68 0.04 0.11 0.95 0.63
3 41.41 15.58 87.28 63.60 22.84 2.82 3.11 1.05
4 14.34 2.34 56.61 34.81 1.66 10.87 10.39 0.48
GCA Tkl Blocks Int Clr Err
1 0.09 0.36 0.06 0.26 6.87 1.34
2 0.59 39.94 34.80 34.11 119.34 1.01
3 8.64 56.30 29.47 24.78 25.48 0.43
4 7.48 7.45 8.20 2.13 6.13 0.10
The profile means can be interpreted in the following way:
  • Profile 1 consistently exhibits low values compared to the other profiles across almost every variable, except for Err – Errors. This profile can be interpreted as grouping the goalkeepers, as they typically have significantly lower average values in the considered variables. These variables are more relevant to players in more active roles involving ball possession.
  • Profile 2 is characterized by notably high values in Clr – Clearances, Int – Interceptions, Blocks – Number of times blocking the ball by standing in its path, and Tkl – Tackles. Additionally, it exhibits relatively high values in Err – Errors and Def – Defensive actions that lead to a shot attempt compared to the other profiles. These characteristics are commonly associated with defenders.
  • Profile 3 displays remarkably high values in Tkl – Tackles, GCA – Goal-Creating Actions, Def – Defensive actions that lead to a shot attempt, PassDead – Completed dead-ball passes that lead to a shot attempt, PassLive – Completed live-ball passes that lead to a shot attempt, SCA – Shot-Creating Actions, CrsPA – Crosses into Penalty Area, PPA – Passes into Penalty Area, PiFT – Passes into Final Third, KP – Key Passes, PrgP – Progressive Passes, PrgC – Progressive carries. Additionally, it exhibits high values in Ast – Number of assists. These characteristics, combined with the consistently high values across all these variables, suggest an association with full-backs and midfielders. These positions involve moving the ball around the field and are frequently positioned centrally during the game.
  • Profile 4 demonstrates remarkably high values in ShToSh – Shots that lead to another shot attempt, TO – Successful take-ons that lead to a shot attempt, SoT – Shots on Target, Sh – Shots Total, and Gls – Number of goals. These characteristics are typically associated with pure strikers.
Figure 3 represents the memberships of the archetypal players (very high membership score) for each profile and a selection of other players that exhibit an interesting profile memberships. The archetypal players are:
  • Goalkeepers: Luigi Sepe of Salernitana team, who has a membership 0.962 in Profile 1.
  • Defenders: Federico Baschirotto of Lecce, who has a membership 0.973 in Profile 2.
  • Full-backs and midfielders: Cristiano Biraghi of Fiorentina, who has a membership 0.963 in Profile 3.
  • Pure strikers: Victor Osimhen of Napoli, who has a membership 0.947 in Profile 4.
Among the other players displayed, Chris Smalling, primarily a defensive player for Roma, but he also demonstrates notable scoring ability, having netted several goals in the season under review. This dual aspect of his play is captured by our model, which assigns him a modest membership to the pure strikers profile, acknowledging this secondary role. Edin Dzeko’s well-rounded abilities as a complete centre forward, blending technique with game vision, are reflected in his affiliation to the midfielders’ profile. Paulo Dybala, one of Serie A’s most technically skilled and elegant second forwards, often takes on ball-running duties in the forward area. This blend of a striker’s and a midfielder’s responsibilities is captured in our model. Likewise, Henrikh Mkhitaryan, an unconventional and agile trequartista, showcases a hybrid skill set that encompasses both attacking prowess and defensive contributions. Lastly, Wilfried Singo, known for his physical strength and speed as a right full-back or outside midfielder, is characterized by his dynamic play, adept in both offensive drives and defensive recoveries. As illustrated in Figure 3, the partial membership model adeptly encapsulates these diverse player characteristics, demonstrating its efficacy in capturing the complex nature of football players’ roles and abilities.
Figure 3: Partial membership model. A graphical representation of the archetypal football players and a selection of other interesting players. The pie chart for each player shows their partial membership to the profiles.
Based on the findings presented in Figure 3, it is evident that the model effectively captures the playing positions of the football players and successfully highlights nuances in their playing styles. The profile membership values and pie charts for each player in the database are displayed in the supplementary files of this paper.
These results could have practical implications for coaches, talent scouts, team managers, and analysts. These stakeholders can utilize the findings to make informed decisions related to team strategy, talent acquisition, and statistical research, ultimately enhancing performance and understanding in the field of football.
3.2 Mixed Membership Model Application
Mixed membership models are applied to count data in White and Murphy (2016). A detailed description of the model is given therein.
The profiles means for the mixed membership (MM) model applied to the football player performance data are presented in Table 5 and Figure 4. It must be emphasised, that in this application, the MM model fails in capturing the cited above archetypes. The players with highest membership per profile are: the striker Ciro Immobile with 0.591 in profile 1, the defensive player Gianluca Mancini with 0.485 membership in profile 2, the defensive player Martin Erlic with 0.700 membership in profile 3, the midfielder Jerdy Schouten with 0.533 membership in profile 4 and the second-tier Christian Kouamé with 0.548 membership in profile 5.
@@ 3.2 Mixed Membership Model Application
Table 5: Mixed membership model estimated expected profile means.
Gls Ast PrgC PrgP Sh SoT KP PiFT
1 7.34 3.91 64.14 26.44 35.62 16.83 20.35 109.46
2 1.25 1.04 19.66 73.55 55.81 5.89 14.09 10.32
3 0.12 0.11 7.86 0.19 0.25 0.31 0.93 42.31
4 0.74 0.90 0.13 116.27 9.17 2.81 6.34 66.60
5 2.35 2.36 37.30 48.66 20.00 7.80 37.36 25.83
PPA CrsPA SCA PassLive PassDead TO ShToSh Def
1 31.13 1.60 44.59 1.21 0.45 5.79 5.90 0.50
2 10.64 3.68 30.02 22.25 20.73 0.98 1.76 0.72
3 1.01 0.11 10.24 8.11 0.52 0.04 0.17 0.17
4 4.77 1.64 1.99 32.39 1.09 0.36 0.93 0.76
5 18.92 11.59 72.56 49.05 5.04 2.25 3.12 0.96
GCA Tkl Blocks Int Clr Err
1 8.77 10.51 10.01 4.51 6.17 0.22
2 2.42 0.96 17.60 29.90 16.28 0.33
3 0.35 22.30 0.10 17.71 84.78 1.17
4 2.35 33.45 31.19 15.49 20.77 0.64
5 5.28 50.20 19.57 0.36 46.82 0.27
Figure 4: Mixed membership model. Expected profiles means, conditional on profile membership, with 5 profiles.
The interpretation of the profile means in the MM model is far to be immediate compared to the one of PM model.
  • Profile 1 suggests a grouping of strikers, but it also exhibits high mean values in variables more typical of midfielders, such as PrgC – Progressive carries.
  • Profile 2 shows high mean values in Int – Interceptions, Blocks – Number of times blocking the ball by standing in its path, PassDead – Completed dead-ball passes that lead to a shot attempt, and Sh – Shots Total. These characteristics could be associated with defensive midfielders, although Sh - Shots Total is not typical for this role.
  • Profile 3 generally has low values in most variables, indicating a potential grouping of goalkeepers. However, it also presents high values in Int — Interceptions and Tkl – Tackles, which are more commonly associated with defenders.
  • Profile 4 exhibits very high values in PrgP – Progressive Passes, Blocks – Number of times blocking the ball by standing in its path, and Tkl – Tackles. This profile could be interpreted as grouping offensive midfielders, even though Blocks are not typically associated with this role.
  • Profile 5 might group defenders and full-backs, but these two positions have distinct characteristics, making the interpretation somewhat challenging.
The membership values resulted by the MM model and the respective pie charts for each player in the database are displayed in the supplementary files of this paper.
3.3 Finite Mixture Model Application

### v2

@@ 2.6 Model Selection
According to Watanabe and Opper (2010), a statistical model is said to be regular if the map taking parameters to probability distributions is one-to-one and if its Fisher information matrix is positive definite. If a model is not regular, then it is said to be singular. If a statistical model contains a hierarchical structure or latent variables, then the model is generally singular. In singular statistical models, the maximum likelihood estimator does not satisfy asymptotic normality. Consequently, standard model selection criteria like the Akaike Information Criterion (AIC) and the Bayes Information Criterion (BIC) are not always appropriate for model selection.
WAIC (Widely Applicable Information Criterion) (Watanabe and Opper, 2010) can be used for estimating the predictive loss of Bayesian models, using a sample from the full-data posterior, and it is applicable to non-regular models such as the PM. In the present paper, WAIC is calculated from Equations 5, 12, and 13 in Gelman et al (2014), and it is the log pointwise predictive density minus a correction for effective number of parameters to adjust for overfitting. According to Millar (2018), the marginalized WAIC might be more accurate for choosing the correct model. Indeed, in our simulation (detailed below), we compare the WAIC conditional on all of the parameters (WAICc) with the WAIC marginalized for $\bm{\tau}$ (WAICm) and find that WAICm more reliably detects the true $K$ in a partial membership setting.
In our real-data application, we report the $\text{WAIC}_{m}$. In mixture model application it has been marginalized over the mixing proportions $\bm{\pi}$.
We fit the model in a simulated data scenario to verify the number of times the correct model is selected by the WAIC conditional on all of the parameters ($\text{WAIC}_{c}$) and marginalized over $\bm{\tau}$ ($\text{WAIC}_{m}$). We generated 100 random membership vectors from a Dirichlet($\bm{\delta}$) distribution with shape parameter $\bm{\delta}=(0.5,0.5,0.5,0.5)$. Using these membership scores, we simulated 100 partial membership models with $N=200$, $J=25$, and $K=4$ to match the football players application scenario. Each $\lambda_{kj}$ was drawn from an $\mbox{Exponential}(1/5)$ distribution, yielding an average rate of 0.2. Additionally, each $\lambda_{kj}$ in the partial membership model was assigned a $\mathrm{Gamma}(1,1)$ prior, as this relatively uninformative choice had little effect on the estimated clusters. We ran the MCMC algorithm for 20000 iterations, keeping every 50th draw, and discarded the first 5000 iterations as burn-in. MCMC was run for $K=1,\dots,8$, and convergence was assessed by examining trace plots.
The simulation results are illustrated in Table 2, confirming that the marginalized WAIC ($\text{WAIC}_{m}$) is more accurate for choosing the correct model, with 99% of runs favoring the true $K=4$.
Table 2: Number of times each $K$ yields the minimum WAIC conditional on parameters ($\text{WAIC}_{c}$) and marginalized over $\bm{\tau}$ ($\text{WAIC}_{m}$). $N=200$, $J=25$, Number of runs=100, true $K=4$. (Values in bold indicate the best-performing $K$.)
$\text{WAIC}_{c}$ $\text{WAIC}_{m}$
@@
Finally, we note that WAIC values from partial membership, mixed membership, and finite mixture models may appear on different numeric scales because each model specifies a distinct likelihood factorization. In Table 3, for example, the ($\text{WAIC}_{m}$) values for PM can be somewhat larger than for MM or mixture. However, the primary goal is to compare models within each class for a given dataset and to note which model achieves the lowest WAIC overall.
@@ 3 Serie A Football Players
We analyze the performance statistics of 200 Serie A football players who played more than 1720 minutes during the 2022/2023 football season [^1] The data are freely available at https://fbref.com/en/comps/11/2022-2023/stats/2022-2023-Serie-A-Stats.. The analysis considers a set of 22 count variables for each player across the games that they played, selected to encompass the essential skills associated with each player’s role on the field.
@@
The performance variables included in the analysis and modeling are given as follows.
@@
Partial membership (PM), mixed membership (MM) and mixture models are applied to the dataset, with the $\text{WAIC}_{m}$ suggesting that the $K=4$, $K=5$ and $K=6$ profile models are optimal, as illustrated in Table 3. All three models are estimated under a Bayesian framework using the same general MCMC procedure described in Section 2.6. Specifically, we implement each model in Nimble, specify priors for all parameters, run a burn-in and thinning period, and then derive posterior summaries for inference. Although we summarize results (such as membership vectors or cluster parameters) via posterior means and highlight the most likely cluster assignments, these point estimates remain rooted in the Bayesian posterior. This ensures that all uncertainty quantification and model selection (based on $\text{WAIC}_{m}$) are conducted consistently across the three model types, while any “frequentist-looking” statistics (like means or modes) are simply convenient summaries of the posterior distribution.
Table 3: $\text{WAIC}_{m}$ values (rounded to the nearest integer) for the partial membership (PM), mixed membership (MM), and mixture models, with $K$ ranging between 2 and 8. The optimal model for each model type is in bold.
$K=2$ $K=3$ $K=4$ $K=5$
PM 46137 36360 33415 37796
MM 26823 22702 21330 20860
Mixture 30816 24648 23090 21948
@@ 3.1 Partial Membership Model Results
Table 4 and Figure 2 display the posterior mean of each profile’s Poisson parameters for the partial membership (PM) model with $K=4$. The PM model allows each player to lie anywhere in the simplex of these four profiles, thus naturally capturing hybrid or multi-role behaviors.
@@
Table 4: PM model expected profiles means (Poisson rate parameters).
Gls Ast PrgC Sh KP CrsPA SCA TO
1 20.91 6.26 58.57 111.62 49.21 3.64 99.73 12.50
2 2.44 3.90 103.65 32.33 51.82 24.35 98.56 3.56
3 1.32 0.34 6.84 14.68 3.43 0.25 13.19 0.16
4 0.00 0.09 0.77 0.01 0.25 0.15 2.38 0.02
ShToSh Def GCA Tkl Blocks Int Clr
1 12.24 0.72 13.71 11.59 10.91 3.58 5.40
2 3.71 1.82 8.90 89.44 45.86 42.11 49.27
3 1.79 0.60 1.03 47.97 41.96 40.73 133.32
4 0.00 0.07 0.22 0.51 0.25 0.36 9.07
The four profiles can be interpreted as follows:
  • Profile 1: Exhibits exceptionally high values in Shots (Sh), Goals (Gls), Assists (Ast), Key Passes (KP), and Shot-Creating Actions (SCA). The notable presence of Take-Ons (TO) and Shots that lead to another shot attempt (ShToSh) further underscores an aggressive, forward-oriented playing style. This profile aligns with strikers or highly offensive forwards who focus on scoring and creating chances.
  • Profile 2: Remarkably high in Progressive Carries (PrgC), Key Passes (KP), Crosses into the Penalty Area (CrsPA), and Shot-Creating Actions (SCA), yet also excels defensively with large counts of Tackles (Tkl), Blocks, and Interceptions (Int). The presence of moderate Goals (Gls) and Assists (Ast) suggests a well-rounded role, typical of full-backs and dynamic midfielders who contribute in both offensive and defensive phases.
  • Profile 3: Dominated by very high Clearances (Clr), with substantial Tackles, Blocks, and Interceptions as well. Offensively, it registers relatively low values in Goals and Shots, suggesting that these players rarely assume forward duties. These characteristics match classical defenders, especially center-backs, who prioritize marking, blocking, and clearing the ball.
  • Profile 4: Displays consistently low counts for nearly all offensive and defensive variables, aside from a modest number of Clearances (Clr). This profile is best interpreted as goalkeepers, who typically do not generate large numbers of shots, passes, or tackles. Clearances may occur when the keeper acts as a sweeper or clears loose balls, but otherwise these players’ statistics remain low across the board.
The estimated memberships $\bm{\tau}_{i}$ are represented in the tetrahedron in Figure 3. This plot provides a view of how each player’s profile splits across the four clusters: players near a corner (e.g., Victor Osimhen, Sebastiano Luperto, Rogério and Alex Meret) show almost “pure” membership in one cluster, whereas those scattered within the interior exhibit genuine multi-role behaviors.
Figure 3: Tetrahedron Representation of player profiles memberships.
Figure 4 illustrates the memberships of the archetypal players and other interesting players who split their membership. Notably:
  • Victor Osimhen (Napoli) is heavily associated with Profile 1 with membership $0.904$, confirming his striker role.
  • Rogério Oliveira da Silva (Sassuolo) has strong membership in Profile 2, namely $0.926$, aligning with his full-back/midfielder responsibilities.
  • Sebastiano Luperto (Empoli) fits almost exclusively into Profile 3 with $0.984$ membership, exemplifying a traditional defender’s style.
  • Alex Meret (Napoli) anchors Profile 4 with $0.962$ membership, as a classic goalkeeper.
Additional players such as Paulo Dybala and Henrikh Mkhitaryan demonstrate partial memberships reflecting hybrid roles that draw from both offensive and midfield profiles. Nicolò Barella’s split between Profiles 1 and 2 illustrates a box-to-box midfielder with notable attacking flair, while Marcelo Brozovic’s combination of Profiles 2, 1, and 3 underscores his deep-lying playmaking and defensive screening abilities. Lastly, Wilfried Singo, known for his physical strength and speed as a right full-back or outside midfielder, is characterized by his dynamic play, adept in both offensive drives and defensive recoveries. As illustrated in Figure 4, the partial membership model encapsulates these diverse characteristics of the players well, demonstrating its efficacy in capturing the complex nature of the roles and abilities of football players. The expected values of $\tau$, estimated as posterior means of the partial membership weights, represented for nine players in Figure 4, summarize the relative contributions of each cluster to a player’s overall profile. For example, a player with $\tau=(0.25,0.25,0.25,0.25)$ is interpreted as balancing characteristics between four roles, rather than belonging exclusively to one with uncertainty. The resulting insights can be valuable for coaches, scouts, and analysts who wish to understand and leverage the multi-faceted contributions of each player on the pitch.
Figure 4: Partial membership model. Selected players’ membership vectors, with pie charts illustrating their fractional membership in each of the 4 profiles.
3.2 Mixed Membership Model Results
Mixed membership (MM) models for count data are discussed in White and Murphy (2016), and a detailed description of their structure is given therein. We fit an MM model to the same Serie A player data under consideration. The estimated Poisson means for each of the five profiles are shown in Table 5 and visualized in Figure 5.
Although the MM framework allows an observation’s individual attributes (e.g. passes, tackles) to be drawn from different clusters, it does not identify “archetypal” players who fully belong to a single role (as partial membership models might). In practice, this means the MM model may capture hybrid behaviors, but interpreting each profile can be less straightforward than with partial membership.
@@ 3.2 Mixed Membership Model Results
Table 5: Mixed membership model: expected profile means (Poisson rate parameters).
Gls Ast PrgC Sh KP CrsPA SCA TO
1 2.68 1.89 45.78 0.09 27.33 18.70 58.49 2.90
2 0.92 1.04 21.96 19.31 14.96 2.22 24.37 0.39
3 2.88 3.49 82.51 8.89 5.15 7.58 96.11 2.58
4 10.13 4.95 0.05 62.69 0.76 1.96 42.17 7.83
5 0.16 9.70 34.11 51.36 0.17 7.65 0.06 0.23
ShToSh Def GCA Tkl Blocks Int Clr
1 3.13 1.07 5.25 27.41 41.71 9.52 57.39
2 1.27 1.09 2.53 67.04 0.14 24.05 16.96
3 4.01 1.34 6.84 42.14 27.60 40.21 29.79
4 7.36 0.44 11.53 12.05 12.74 1.60 7.18
5 0.23 0.61 1.14 21.00 18.85 98.51 0.05
Figure 5: MM model. Expected profile means, conditional on profile membership.
We now provide a tentative interpretation of each profile based on the most prominent variables:
  • Profile 1. Moderately high progressive carries (PrgC) and notably large values for shot-creating actions (SCA) and clearances (Clr). The combination of moderate offensive attributes (KP, Ast) with strong defensive indicators (Blocks, Tkl, Clr) could represent wide defenders or full-backs who frequently move the ball forward and also clear threats.
  • Profile 2. More balanced attacking contributions (Sh $=19.31$, KP $=14.96$) but a standout tackle count (Tkl $=67.04$) and significant interceptions (Int $=24.05$). This profile suggests a defensive-minded midfielder, someone who can take a moderate number of shots but excels at reclaiming possession and breaking up play.
  • Profile 3. Very high progressive carries (82.51) and SCA (96.11), combined with substantial defensive metrics (Tkl $=42.14$, Blocks $=27.60$, Int $=40.21$). This cluster may reflect all-around midfielders or possibly wide players who contribute heavily to both ball progression and defense, to box-to-box midfielders or very active wing-backs.
  • Profile 4. Strikingly high goals (10.13), shots (62.69), and shot-to-shot passes (ShToSh $=7.36$), but minimal progressive carries (0.05). This points to a pure striker or center-forward role, someone who focuses on finishing rather than ball-carrying.
  • Profile 5. Low scoring (Gls $=0.16$) but a large number of assists (9.70), many shots (51.36), and an extremely high interceptions count (Int $=98.51$). This combination might characterize a deep-lying playmaker or attacking midfielder who also excels at intercepting passes, suggesting an unusual hybrid of offensive and defensive ability.
Although these labels offer a rough guide, the mixed membership model essentially spreads each player’s attributes across multiple latent profiles. Consequently, each of these profiles may mix features of different roles, making them more challenging to interpret than profiles from the partial membership model. Full membership values for each player are provided in the supplementary materials.
3.3 Finite Mixture Model Results